# ADAPTIVE-$\lambda$ SISS in MACHINE UNLEARNING for DDPM & VAE

### v1

@@ 5.3 Unlearning as a Sequential Decision Process: Reinforcement Learning for $\lambda$ Selection
Traditional unlearning methods, including the Static-$\lambda$ SISS framework and simple heuristic scheduling, primarily employ a greedy approach to loss balancing. These methods optimize $\lambda$ based only on the immediate objective, failing to account for the crucial long-term effects that a single regularization weight choice has on the unlearning stability and the final state of the model. To overcome this fundamental limitation and learn a policy that maximizes long-term, discounted unlearning success, we re-frame the selection of the regularization weight $\lambda_{t}$ at step $t$ as a formal Markov Decision Process (MDP).
5.3.1 MDP Formulation for Adaptive $\lambda$
The unlearning process is modeled as an MDP defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma)$, where the policy $\pi$ learned by the agent determines the optimal $\lambda_{t}$ given the current state.
The State Space ($\mathcal{S}$) is designed to provide the agent with a complete, low-dimensional view of the current unlearning dynamics and history. The state $S_{t}$ at time $t$ captures the instantaneous loss values, the magnitude of the model’s response to these losses via gradient norms, the immediate past action, and the progression of time:
[…]
The Action Space ($\mathcal{A}$) is a single-dimensional continuous space, representing the chosen regularization weight $\lambda_{t}$ for the current step:
$A_{t}=\Big\{\lambda_{t}\in[\lambda_{\min},\lambda_{\max}]\Big\}$
The formulation of the Reward Function ($\mathcal{R}$) is critical, as it encodes the desired trade-off. The immediate reward $R_{t}$ is designed to maximize unlearning efficacy ($\mathcal{L}_{\text{forget}}-\mathcal{L}_{\text{retain}}$), while simultaneously applying penalties to enforce training stability and model safety. The term $\gamma|\lambda_{t}-\lambda_{t-1}|$ penalizes large, erratic changes in the action, thus promoting a smoother unlearning trajectory. The term $\mathcal{L}_{\text{retain}}^{2}$ acts as a strong safety penalty, ensuring the policy prioritizes minimal degradation on the retain set:
$R_{t}=\underbrace{-\mathcal{L}_{\text{retain}}+\mathcal{L}_{\text{forget}}}_{\text{Unlearning Efficacy}}-\underbrace{\gamma|\lambda_{t}-\lambda_{t-1}|}_{\text{Stability Penalty}}-\underbrace{\mathcal{L}_{\text{retain}}^{2}}_{\text{Safety Penalty}}$
5.3.2 RL Algorithms for Policy Optimization
To learn the optimal continuous policy $\pi(\lambda_{t}|S_{t})$, we consider several robust policy optimization algorithms suited for continuous control tasks.
Monte Carlo Policy Gradient (MCPG)
As a foundational method for optimizing a stochastic policy $\pi_{\theta}(\lambda_{t}|S_{t})$, the Monte Carlo Policy Gradient aims to adjust the policy parameters $\theta$ in the direction of higher return. Given the stochastic nature of the $\lambda$ selection (similar to our proposed $\lambda$-encoder from the Variational Inference section), a policy loss can be derived that incorporates a regularization term:
$\mathcal{L}_{\text{policy}}=-R_{t}\log(\pi_{\theta}(\lambda_{t}|S_{t}))+\beta D_{\text{KL}}(q_{\phi}(z)||p(z))$
Here, $R_{t}$ acts as an estimated advantage or return, and the KL divergence term helps to constrain the policy distribution.
Proximal Policy Optimization (PPO)
Proximal Policy Optimization is a highly stable and sample-efficient on-policy algorithm used to perform the optimization. PPO improves upon standard Policy Gradient methods by employing a clipped surrogate objective. This clipping mechanism prevents extreme policy shifts by bounding the probability ratio $r_{t}(\theta)=\frac{\pi_{\theta}(A_{t}|S_{t})}{\pi_{\theta_{\text{old}}}(A_{t}|S_{t})}$. This stability is crucial in the volatile setting of machine unlearning, as it prevents the policy from converging to poor local optima early in the training process. The clipped objective is designed to maximize the following expression, where the clipping region prevents drastic divergence:
$\max_{\theta}\mathbb{E}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$
Deep Deterministic Policy Gradient (DDPG) with Entropy Regularization
DDPG is an established off-policy algorithm for deterministic continuous control, where the policy directly outputs the continuous $\lambda_{t}=\pi_{\theta}(S_{t})$. To ensure adequate exploration in the complex $\lambda$ space, especially early in the process, the deterministic action is often perturbed by noise: $\lambda_{t}=\pi_{\theta}(S_{t})+\epsilon$. Furthermore, the algorithm can be modified to incorporate maximum entropy regularization, similar to methods like Soft Actor-Critic (SAC). The objective function is augmented with the policy’s entropy $\mathcal{H}(\pi)$ to encourage broader exploration:
$\mathcal{L}=\mathbb{E}[Q(S,A)]+\alpha\mathcal{H}(\pi)$
While DDPG is considered the "Best Continuous Control" framework for many applications, the use of a fully entropy-regularized approach for a single-dimensional action space $\lambda$ may introduce unnecessary complexity. However, the benefits of improved exploration and robustness could still outweigh the implementation overhead.

### v2

@@ 5.3 Unlearning as a Sequential Decision Process: Reinforcement Learning for $\lambda$ Selection
The adaptive SISS formulation in Section 5.1 makes $\lambda$ depend on the current training context, but the optimization is still essentially myopic: at each step, $\lambda$ is chosen to reduce the instantaneous SISS loss on the current mini-batch. In practice, however, a particular choice of $\lambda$ can have long-term effects on the model. Choosing an overly aggressive forgetting weight early in training can irreversibly damage the retain distribution, while overly conservative values may lead to under-forgetting even after many optimization steps.
To explicitly capture this long-term trade-off between forgetting and retention, we model the unlearning process as a finite-horizon Markov Decision Process (MDP), in which adaptively selecting $\lambda$ becomes a sequential decision problem, and the environment dynamics are induced by the gradient updates of the diffusion model.
5.3.1 MDP formulation
We define an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma)$ that describes the evolution of the model during unlearning.
Episodes and time index.
We consider unlearning as a finite-horizon process of $T$ optimization steps. One episode corresponds to a full unlearning run of $T$ gradient updates. At each step $t=1,\dots,T$, an agent observes the current unlearning state of the model, chooses a forgetting weight $\lambda_{t}$, and the model parameters are updated using the SISS loss evaluated with that choice of $\lambda_{t}$.
State space.
The state $S_{t}\in\mathcal{S}$ should summarize the current “unlearning status” of the model. We build on the context features already used in the adaptive SISS framework and augment them with the previous action and a notion of training progress. A simple and effective state representation is
$S_{t}=\big[\mathcal{L}_{\text{retain},t},\,\mathcal{L}_{\text{forget},t},\,\|\nabla_{\theta}\mathcal{L}_{\text{retain},t}\|,\,\|\nabla_{\theta}\mathcal{L}_{\text{forget},t}\|,\,\lambda_{t-1},\,t/T\big],$ (15)
where $L_{\text{retain},t}$ and $L_{\text{forget},t}$ are mini-batch estimates of the retain and forget losses at step $t$, and $\|\nabla_{\theta}\mathcal{L}_{\text{retain},t}\|$, $\|\nabla_{\theta}\mathcal{L}_{\text{forget},t}\|$ are the corresponding gradient norms. This state vector provides a compact, approximately Markovian summary of how the current model responds to the retain and forget objectives.
Action space.
The action $A_{t}\in\mathcal{A}$ is the choice of the forgetting weight $\lambda_{t}$ at step $t$:
$\mathcal{A}=\big\{\lambda\in[\lambda_{\min},\lambda_{\max}]\big\}\subset\mathbb{R},$ (16)
where $\lambda_{\min},\lambda_{\max}]$ is a predefined admissible range. In practice, this range can be chosen as a strict subset of $[0,1]$ to avoid degenerate cases such as $\lambda=0$ (no forgetting) or $\lambda=1$ (overly aggressive forgetting).
Transition dynamics.
Given a state $S_{t}$ and an action $\lambda_{t}$, the environment evolves as follows:
  1. 1.
    Sample a mini-batch $x_{r}\sim P_{R},\,x_{f}\sim P_{F},\,t^{\prime}\sim P_{T}$, and $m_{t^{\prime}}\sim q_{\lambda_{t}}(m_{t^{\prime}}\mid x_{r},x_{f})$ as in the SISS formulation.
  2. 2.
    Compute the SISS loss $\mathcal{L}_{\text{siss},t}(\lambda_{t})$ for this mini-batch.
  3. 3.
    Update the diffusion model parameters by one gradient step:
    $\theta_{t+1}\leftarrow\theta_{t}-\eta\,\nabla_{\theta}\mathcal{L}_{\text{siss},t}(\lambda_{t}),$ (17)
    where $\eta$ is the learning rate.
  4. 4.
    Recompute the losses and (optionally) their gradient norms under $\theta_{t+1}$ to form the next state $S_{t+1}$.
Because of mini-batch sampling and the stochastic mixture $q_{\lambda_{t}}(m_{t^{\prime}}\mid x_{r},x_{f})$, the transition dynamics $S_{t}\mapsto S_{t+1}$ are stochastic.
Reward function.
Our ultimate objective is to minimize the cumulative SISS loss over the entire unlearning trajectory, while maintaining stability and protecting the retain distribution. We therefore define a step-wise cost as
$C_{t}(\lambda_{t})=\mathcal{L}_{\text{siss},t}(\lambda_{t})+\alpha\big(\lambda_{t}-\lambda_{t-1}\big)^{2},$ (18)
where $\alpha>0$ penalizes abrupt changes in $\lambda$ between consecutive steps. The corresponding reinforcement learning reward is simply the negative cost:
$R_{t}=r(S_{t},\lambda_{t})=-C_{t}(\lambda_{t})=-\mathcal{L}_{\text{siss},t}(\lambda_{t})-\alpha\big(\lambda_{t}-\lambda_{t-1}\big)^{2}.$ (19)
Under this definition, maximizing the expected discounted return is equivalent to minimizing the expected cumulative SISS loss plus a smoothness regularizer on the sequence of forgetting weights.
Objective.
Let $\pi_{\psi}(\lambda\mid S)$ denote a stochastic policy with parameters $\psi$ that maps states to actions (i.e., selects $\lambda$ based on $S$). Let $\tau=(S_{1},\lambda_{1},S_{2},\lambda_{2},\dots,S_{T},\lambda_{T})$ denote an unlearning trajectory induced by this policy and by the stochastic dynamics of the environment. The reinforcement learning objective is to maximize the expected discounted return
$J(\psi)=\mathbb{E}_{\tau\sim p_{\psi}(\tau)}\bigg[\sum_{t=1}^{T}\gamma^{t-1}R_{t}\bigg]=\mathbb{E}_{\tau\sim p_{\psi}(\tau)}\bigg[-\sum_{t=1}^{T}\gamma^{t-1}\Big(\mathcal{L}_{\text{siss},t}(\lambda_{t})+\alpha(\lambda_{t}-\lambda_{t-1})^{2}\Big)\bigg],$ (20)
where $\gamma\in(0,1]$ is a discount factor. The optimal dynamic $\lambda$-selection strategy is then given by a policy $\pi^{\star}_{\psi}$ that maximizes $J(\psi)$.
5.3.2 Policy parameterization and relation to adaptive $\lambda$
To keep the reinforcement learning formulation consistent with the adaptive SISS framework, we parameterize the policy as a Gaussian latent policy followed by a squashing function that maps the latent variable into the admissible range of $\lambda$.
Given a state $S_{t}$, the policy network $f_{\psi}$ produces the mean and standard deviation of a Gaussian:
$\big(\mu_{\psi}(S_{t}),\,\sigma_{\psi}(S_{t})\big)=f_{\psi}(S_{t}).$ (21)
We then sample a latent variable $z_{t}$ using the reparameterization trick:
$z_{t}=\mu_{\psi}(S_{t})+\xi_{t}\,\sigma_{\psi}(S_{t}),\quad\xi_{t}\sim\mathcal{N}(0,1),$ (22)
and obtain the action $\lambda_{t}$ by squashing $z_{t}$ into the interval $\lambda_{\min},\lambda_{\max}]$:
$\lambda_{t}=\lambda_{\min}+(\lambda_{\max}-\lambda_{\min})\,\sigma(z_{t}),$ (23)
where $\sigma(\cdot)$ is the sigmoid function. This parameterization is identical in form to the variational encoder for $\lambda$ introduced in Section LABEL:sec:siss_adaptive; the key difference lies in the training objective: instead of minimizing a per-step ELBO, we now optimize the expected long-term return $J(\psi)$.
To encourage sufficient exploration and to keep the policy close to a prescribed prior $p(\lambda)$, we additionally include an entropy- or KL-regularization term. A convenient objective is
$\mathcal{L}_{\text{RL}}(\psi)=-J(\psi)+\beta\,\mathbb{E}_{S_{t}}\big[D_{\mathrm{KL}}\big(\pi_{\psi}(\lambda\mid S_{t})\,\big\|\,p(\lambda)\big)\big],$ (24)
with a regularization weight $\beta>0$. This yields an entropy-regularized reinforcement learning objective that is mathematically analogous to the ELBO used in the variational SISS formulation, but now extended over entire trajectories instead of single optimization steps.
In practice, we optimize $\mathcal{L}_{\text{RL}}$ using standard continuous-control policy gradient methods such as Proximal Policy Optimization (PPO). At each iteration, we roll out several unlearning trajectories under the current policy $\pi_{\psi}$, recording states, selected forgetting weights, rewards, and SISS losses. These trajectories are then used to estimate advantages and update $\psi$ with a clipped policy-gradient objective. The diffusion model parameters $\theta$ are updated throughout these rollouts using standard gradient descent on the SISS loss evaluated at the policy-selected values of $\lambda_{t}$. In this way, the policy learns to choose dynamic, state-dependent forgetting weights that optimize the entire unlearning trajectory, rather than only the instantaneous loss on each mini-batch.
5.3.3 Implementation details for PPO-based adaptive $\lambda$
We now describe how we instantiate the MDP formulation of Section 5.3 in practice and how we train the PPO agent that selects the forgetting weights $\lambda_{t}$.
Environment and episodes.
We treat one environment step as one gradient update of the diffusion model parameters $\theta$ on a single mini-batch. An episode corresponds to $T$ such updates, starting from a fixed pretrained checkpoint $\theta_{0}$. Unless otherwise stated, we set $T=200$ environment steps per episode. At the beginning of each episode, we reset $\theta\leftarrow\theta_{0}$ and initialize the previous forgetting weight to a default value $\lambda_{0}$ (we use $\lambda_{0}=0.5$ in all experiments).
Mini-batch construction.
At each environment step $t$, we construct a mini-batch by sampling a set of “forget” examples from $\mathcal{D}_{\text{forget}}$ and a set of “retain” examples from $\mathcal{D}_{\text{retain}}$, using the same sampling strategy as in the original SISS training. We use a batch size of $B$ samples per step (e.g., $B=64$), with a fixed ratio between forget and retain samples (e.g., $B/2$ from each set). Given the current forgetting weight $\lambda_{t}$ selected by the agent, we sample $m_{t}$ from the mixture $q_{\lambda_{t}}(m_{t}\mid x,a)$ and compute the SISS loss $L_{\text{siss},t}(\lambda_{t})$ on this mini-batch.
State, action, and reward.
The state vector $S_{t}$ at step $t$ is constructed from the current mini-batch losses and the previous forgetting weight:
$S_{t}=\big[\mathcal{L}_{\text{retain},t},\,\mathcal{L}_{\text{forget},t},\,\|\nabla_{\theta}\mathcal{L}_{\text{retain},t}\|,\,\|\nabla_{\theta}\mathcal{L}_{\text{forget},t}\|,\,\lambda_{t-1},\,t/T\big],$ (25)
where $\mathcal{L}_{\text{retain},t}$ and $\mathcal{L}_{\text{forget},t}$ are the retain and forget loss components from the SISS objective on the current mini-batch, and the gradient norms are computed with respect to the diffusion model parameters $\theta$. In ablations, we also consider a simplified variant that omits the gradient norms and only uses the scalar losses, which slightly reduces computation.
The action $A_{t}$ is the scalar forgetting weight $\lambda_{t}$ chosen by the policy. We restrict $\lambda_{t}$ to an interval $\lambda_{\min},\lambda_{\max}]$ with $\lambda_{\min}=0.1$ and $\lambda_{\max}=0.9$ to avoid degenerate extremes (no forgetting or excessively aggressive forgetting).
Given the mini-batch SISS loss $\mathcal{L}_{\text{siss},t}(\lambda_{t})$, we define the reward as the negative loss plus a smoothness penalty:
$R_{t}=-L_{\text{siss},t}(\lambda_{t})-\alpha\big(\lambda_{t}-\lambda_{t-1}\big)^{2},$ (26)
with $\alpha>0$ controlling how strongly we penalize large changes in $\lambda$ between consecutive steps. In our experiments we set $\alpha=0.01$. This reward encourages the policy to reduce the SISS loss while adapting $\lambda_{t}$ gradually over time.
Policy and value networks.
Both the policy $\pi_{\psi}(\lambda\mid S)$ and the value function $V_{\phi}(S)$ are implemented as small multi-layer perceptrons (MLPs) that operate on the state vector $S_{t}$. Unless otherwise specified, we use two hidden layers with 64 units each and $\tanh$ activations.
The policy network outputs the mean and log-standard deviation of a univariate Gaussian:
$(\mu_{\psi}(S_{t}),\log\sigma_{\psi}(S_{t}))=f_{\psi}(S_{t}),$ (27)
from which we sample a latent variable $z_{t}$ using the reparameterization trick,
$z_{t}=\mu_{\psi}(S_{t})+\xi_{t}\,\sigma_{\psi}(S_{t}),\quad\xi_{t}\sim\mathcal{N}(0,1).$ (28)
The forgetting weight is then obtained by squashing $z_{t}$ to the admissible range,
$\lambda_{t}=\lambda_{\min}+(\lambda_{\max}-\lambda_{\min})\,\sigma(z_{t}),$ (29)
where $\sigma(\cdot)$ is the sigmoid nonlinearity. The value network $V_{\phi}(S_{t})$ shares the same architecture but outputs a single scalar estimating the expected return from state $S_{t}$.
PPO training.
We optimize the policy and value networks using Proximal Policy Optimization (PPO). For each PPO update, we collect $N$ episodes of length $T$ by interacting with the environment as described above, starting each episode from the same pretrained checkpoint $\theta_{0}$. During data collection, the diffusion model parameters $\theta$ are updated online at every environment step using one gradient step on the SISS loss $\mathcal{L}_{\text{siss},t}(\lambda_{t})$, while the policy parameters $\psi$ and value parameters $\phi$ are kept fixed. The resulting trajectories $\{(S_{t},\lambda_{t},R_{t},S_{t+1})\}$ are stored for PPO.
Given the collected trajectories, we compute discounted returns and generalized advantage estimates (GAE) with discount factor $\gamma=0.99$ and GAE parameter $\lambda_{\mathrm{GAE}}=0.95$. We then perform several epochs of PPO updates on mini-batches of the stored transitions, using the clipped objective with ratio $r_{t}(\psi)=\pi_{\psi}(\lambda_{t}\mid S_{t})/\pi_{\psi_{\text{old}}}(\lambda_{t}\mid S_{t})$ and clipping parameter $\epsilon=0.2$. We use Adam with a learning rate of $3\times 10^{-4}$ for both actor and critic, and include a small entropy bonus to encourage exploration.
After PPO training has converged, we fix the learned policy $\pi_{\psi}(\lambda\mid S)$ and perform a final unlearning run, during which $\theta$ is updated using SISS loss with $\lambda_{t}$ chosen by the trained policy at each step. All unlearning metrics reported in the main text are computed from this final run.
@@ 7 Conclusion
Appendix C: RL Experiment details
We now describe two practical strategies for training the reinforcement learning agent that selects the forgetting weights $\lambda_{t}$ in the MDP formulation of Section 5.3. Both are based on continuous-control algorithms with stochastic policies; in the main experiments we use PPO, and we additionally verify that Soft Actor–Critic (SAC) is also applicable to this setting.
Throughout this section, we treat one environment step as one gradient update of the diffusion model parameters $\theta$ on a single mini-batch, and we use the same state, action, and reward definitions as in Section 5.3. In particular, the state is
$S_{t}=\big[L_{\text{retain},t},\,L_{\text{forget},t},\,\|\nabla_{\theta}L_{\text{retain},t}\|,\,\|\nabla_{\theta}L_{\text{forget},t}\|,\,\lambda_{t-1},\,t/T\big],$ (30)
the action is the scalar forgetting weight $\lambda_{t}\in[\lambda_{\min},\lambda_{\max}]$ with $\lambda_{\min}=0.1$ and $\lambda_{\max}=0.9$, and the reward is
$R_{t}=-L_{\text{siss},t}(\lambda_{t})-\alpha\big(\lambda_{t}-\lambda_{t-1}\big)^{2},$ (31)
with $\alpha=0.01$ unless otherwise noted.
Mini-batch construction.
At each environment step $t$, we construct a mini-batch of size $B$ by sampling from the retain set $\mathcal{D}_{\text{retain}}$ and forget set $\mathcal{D}_{\text{forget}}$. Unless otherwise specified, we use $B=64$ with a balanced split between retain and forget samples. Given the current forgetting weight $\lambda_{t}$ selected by the policy, we sample $m_{t}\sim q_{\lambda_{t}}(m_{t}\mid x,a)$ as in the SISS formulation and compute the mini-batch SISS loss $L_{\text{siss},t}(\lambda_{t})$, which is used both to update $\theta$ and to compute the reward $R_{t}$.
Policy and value architectures.
The policy $\pi_{\psi}(\lambda\mid S)$ and value function $V_{\phi}(S)$ are implemented as small multi-layer perceptrons (MLPs) operating on the state vector $S_{t}$. In all experiments, we use two hidden layers with 64 units each and $\tanh$ activations. The policy network outputs the mean and log-standard deviation of a univariate Gaussian,
$(\mu_{\psi}(S_{t}),\log\sigma_{\psi}(S_{t}))=f_{\psi}(S_{t}),$ (32)
from which we sample a latent variable $z_{t}$ using the reparameterization trick,
$z_{t}=\mu_{\psi}(S_{t})+\xi_{t}\,\sigma_{\psi}(S_{t}),\quad\xi_{t}\sim\mathcal{N}(0,1),$ (33)
and squash it into the admissible range via
$\lambda_{t}=\lambda_{\min}+(\lambda_{\max}-\lambda_{\min})\,\sigma(z_{t}),$ (34)
where $\sigma(\cdot)$ is the sigmoid function. The value network $V_{\phi}(S_{t})$ shares the same architecture but outputs a single scalar.
Option 1: Episodic unlearning with parameter resets.
In the first implementation strategy, we construct a “clean” MDP by resetting the diffusion model to a fixed pretrained checkpoint at the start of every episode. One episode consists of $T$ environment steps, each corresponding to one mini-batch update of $\theta$.
  • Episode initialization. At the beginning of each episode, we reset $\theta\leftarrow\theta_{0}$, where $\theta_{0}$ is the pretrained model before unlearning, and initialize the previous forgetting weight to a default value $\lambda_{0}$ (we use $\lambda_{0}=0.5$). We also perform a single warm-up forward pass to compute the initial losses and build $S_{1}$.
  • Within-episode interaction. For each step $t=1,\dots,T$ within an episode:
    1. 1.
      Given the current state $S_{t}$, the policy samples $\lambda_{t}\sim\pi_{\psi}(\cdot\mid S_{t})$.
    2. 2.
      We construct a mini-batch from $\mathcal{D}_{\text{retain}}$ and $\mathcal{D}_{\text{forget}}$, sample $m_{t}\sim q_{\lambda_{t}}(m_{t}\mid x,a)$, and compute the SISS loss $L_{\text{siss},t}(\lambda_{t})$ on this mini-batch.
    3. 3.
      We update the diffusion model parameters by one gradient step:
      $\theta_{t+1}\leftarrow\theta_{t}-\eta\,\nabla_{\theta}L_{\text{siss},t}(\lambda_{t}),$ (35)
      using Adam with learning rate $\eta$ (e.g., $\eta=10^{-4}$).
    4. 4.
      We compute the reward $R_{t}$ from the SISS loss and smoothness penalty, and rebuild the next state $S_{t+1}$ by recomputing $L_{\text{retain},t+1}$ and $L_{\text{forget},t+1}$ (and their gradient norms) under $\theta_{t+1}$.
    5. 5.
      We store the transition $(S_{t},\lambda_{t},R_{t},S_{t+1})$ as a sample for PPO.
  • PPO updates. After collecting $N$ such episodes (each of length $T$), we perform PPO updates on the policy and value networks. We compute discounted returns and generalized advantage estimates (GAE) with discount factor $\gamma=0.99$ and GAE parameter $\lambda_{\mathrm{GAE}}=0.95$, and optimize the clipped PPO objective with clipping parameter $\epsilon=0.2$. We use Adam with learning rate $3\times 10^{-4}$ for both actor and critic and add a small entropy bonus to encourage exploration. The unlearning model parameters $\theta$ used during data collection are discarded at the end of each episode, and do not directly carry over across episodes.
In this option, the unlearning process is simulated repeatedly from the same starting point $\theta_{0}$ in order to train the policy. After convergence of PPO, we fix $\pi_{\psi}$ and perform a final unlearning run starting from $\theta_{0}$, during which $\theta$ is updated using SISS with $\lambda_{t}$ chosen by the trained policy at each step. All reported unlearning metrics for Option 1 are computed from this final run.
Option 2: Online unlearning in a single long run.
The second strategy uses a single, long unlearning trajectory, without resetting the diffusion model parameters between episodes. This option is closer to the practical deployment setting, and we use it as the primary setup in our main experiments.
  • Single long trajectory. We start from the pretrained model $\theta_{0}$ and run unlearning for $T_{\text{total}}$ environment steps (e.g., $T_{\text{total}}\in[10^{3},10^{4}]$ depending on the dataset). At each step $t$, we:
    1. 1.
      Build the state $S_{t}$ from the current losses and previous forgetting weight.
    2. 2.
      Sample an action $\lambda_{t}\sim\pi_{\psi}(\cdot\mid S_{t})$.
    3. 3.
      Sample a mini-batch and compute $L_{\text{siss},t}(\lambda_{t})$.
    4. 4.
      Update $\theta$ by one gradient step on $L_{\text{siss},t}(\lambda_{t})$.
    5. 5.
      Compute the reward $R_{t}$ and the next state $S_{t+1}$.
    6. 6.
      Store the transition $(S_{t},\lambda_{t},R_{t},S_{t+1})$ in a buffer of recent interactions.
  • Periodic PPO updates. Every $M$ environment steps (e.g., $M=200$), we pause data collection and perform PPO updates using the most recent window of transitions (e.g., the last $N_{\text{steps}}=N\cdot T$ steps, matching the effective batch size used in Option 1). We reuse the same hyperparameters as in Option 1: $\gamma=0.99$, $\lambda_{\mathrm{GAE}}=0.95$, $\epsilon=0.2$, and Adam with learning rate $3\times 10^{-4}$.
In this option, the diffusion model parameters $\theta$ continuously evolve over time, and the same trajectory is used both to train the policy and to perform actual unlearning. Despite the non-stationarity induced by the changing model, PPO remains stable in practice and converges to a policy that selects meaningful dynamic forgetting weights. In all main experiments, we start from this online Option 2 and additionally report results for Option 1 as an ablation to isolate the effect of the episodic resets.
Using Soft Actor–Critic (SAC).
Because the action space for $\lambda_{t}$ is continuous and low-dimensional, off-policy actor–critic methods designed for continuous control are also applicable. In particular, we found that Soft Actor–Critic (SAC) can be used as a drop-in replacement for PPO in both Option 1 and Option 2, with minor modifications:
  • The policy $\pi_{\psi}(\lambda\mid S)$ retains the same squashed Gaussian parameterization as above, but is now trained to maximize both expected return and an entropy term, following the SAC objective.
  • We introduce two Q-functions $Q_{\theta_{1}}(S,\lambda)$ and $Q_{\theta_{2}}(S,\lambda)$ implemented as MLPs with the same architecture as the value network, and train them with Bellman backups on transitions $(S_{t},\lambda_{t},R_{t},S_{t+1})$ stored in a replay buffer.
  • The actor is updated by minimizing the SAC policy loss
    $\mathcal{L}_{\text{actor}}(\psi)=\mathbb{E}_{S_{t}\sim\mathcal{D},\,\lambda_{t}\sim\pi_{\psi}}\big[\alpha_{\text{ent}}\log\pi_{\psi}(\lambda_{t}\mid S_{t})-\min_{i=1,2}Q_{\theta_{i}}(S_{t},\lambda_{t})\big],$ (36)
    where $\alpha_{\text{ent}}$ is the entropy temperature and $\mathcal{D}$ denotes the replay buffer.
  • In Option 1, the replay buffer is filled by repeatedly simulating episodes from the reset model $\theta_{0}$; in Option 2, the buffer is populated online during the single long unlearning run.
Empirically, SAC achieves similar qualitative behavior to PPO in terms of learning dynamic, state-dependent forgetting weights, and sometimes converges in fewer environment steps due to its off-policy nature and replay buffer. However, PPO is simpler to integrate into our existing training loop and is used as the default algorithm in the main results, while SAC is included as an alternative in ablation studies.